How would you slove 3 3/5 divided by 2 1/2=
Thanks

Answer by mathwhiz(13)

Step 1: Change them first to improper fractions.
3 3/5 = 18/5
2 1/2 = 5/2
Step 2: Then find their LCM and convert them.
18/5 = 36/10
5/2 = 25/10
Step 3: Find the reciprocal of 25/10.
reciprocal of 25/10 = 10/25
Step 4: Multiply the two fractions.
36/10 * 10/25 = 360/250
Step 5: Reduce it to lowest terms
360/250 = 36/25 or 1 11/25